Section: 1 of 1 Question: 3 of 7 Marks forthls Question: 10 a



Upload answer sheets



An analysis of the data of the partially destroyed dairy are



variance x=8 and the least square regression lines for the



random variables are 3x 4y = 20, 12x — 8y = 62. Find



the following basis of above infonnation:



(i) Mean value of x and y.



(ii) Correlation co-efficient between x and y.



(iii) SD of x and y.



Ortce you upload f'!es trom your second device, ciid on

1
Expert's answer
2021-12-15T17:31:27-0500

"i)"

Since two regression lines always intersect at a point "(\bar{x},\bar{y})", representing mean values of the values x and y as shown below:

"3x+4y=20"

"12x-8y=62"

Multiplying the first equation by 4 and subtracting from the second,

"24y=18\implies\bar{y}=0.75"

From the first equation

"\bar{x}={(20-4\bar{y})\over3}={17\over3}=5.67"

Therefore, the mean value of x and y are 5.67 and 0.75 respectively.


"ii)"

To find the given regression equations in such a way that the coefficient of dependent variable is less than one at least in one equation. 

So, "3x+4y=20\implies4y=20-3x\implies y=5-{3\over4}x"

That is, "|m_{yx}|=|-{3\over4}|={3\over4}"

and

"12x-8y=62\implies12x=62+8y\implies x={62\over12}+{8\over12}y"

Thus, "m_{xy}={8\over 12}"

Hence coefficient of correlation "r" between "x" and "y" is given by: 

"r=\sqrt{m_{xy}*m_{yx}}=\sqrt{{3\over4}*{8\over12}}=\sqrt{1\over2}=0.7071(4dp)"  


"iii)"

 Since the variance of x is 8, therefore, its standard deviation is,

"\sigma_x=\sqrt{8}=2.8284(4dp)"

To determine the standard deviation of y, consider the formula

"\sigma_y={m_{yx}\sigma_x\over r}=({0.75*2.8284\over0.7071})=3"

Therefore, the standard deviation for x and y are 2.8284 and 3 respectively.